OpenStudy (anonymous):

The three sides of a triangle are 5 cm, 6 cm, and 10 cm. What is the best description for this triangle?

OpenStudy (cwrw238):

square all sides of the triangle and compare the value for the longest sides with the values of the other 2 sides

OpenStudy (cwrw238):

* longest side

OpenStudy (cwrw238):

10^2 = 100 5^2 = 25 6^2 = 36

OpenStudy (cwrw238):

if, as in this case, c^2 > a^2 + b^2 where c is the longest side of the triangle , then the triangle is an obtuse angled triangle

OpenStudy (cwrw238):

if c^2 = a^2 + b^2 it is right-angled if c^2 < a^2 + b^2 it is an acute-angled triangle